# Inflated hot Jupiters: inferring average atmospheric velocity via Ohmic models coupled with internal dynamo evolution

### v1

@@ B.1 Jovian magnetic field curvature
where we have separated the internal (dynamo-generated) and external (magnetosphere/magneto-disk) components. Public data is given in terms of the set of constants $g_{n}^{m}$, $h_{n}^{m}$, $G_{n}^{m}$ and $H_{n}^{m}$ (the Schmidt coefficients) which define the spherical harmonic expansions up to some specific multipole degree using the Schmidt quasi-normalized associated Legendre polynomials $P_{n}^{m}(\cos\theta)$,
@@
Figure 12: Inverse of the mean curvature modulus for different Jupiter magnetic field models as a function of radius, obtained using a discretized version of equation (26). The JRM33 models do not significantly change if the external multipoles are included.
In Fig. 12, we show the inverse of the mean curvature at different radii for both models, including two different cut-offs for JRM33. The gray zone approximately marks the possible start of the conductive zone where the gradual appearance of metallic hydrogen would prevent the validity of the potential solution. The inverse curvature has a mean value of the order of $1/\kappa\sim 0.1\leavevmode\nobreak\ R_{j}$ at $r\sim(0.8-0.9)\leavevmode\nobreak\ R_{j}$ (where the dynamo surface should be), which can be taken as a lower limit of the magnetic field curvature inside the Jovian dynamo.
@@ B.2 Details of MHD dynamo simulations
For example, for the models shown in Fig. 9, the background MESA profiles correspond to an irradiated $1M_{j}$ planet with internal Ohmic heating at 10 Gyr. At that time, the simulated planetary radius is $1.218\leavevmode\nobreak\ R_{j}$, corresponding to an external pressure of and 64 and 5.4 kbar, respectively. We used resolutions of $(N_{r},N_{\theta},N_{\phi})=(193,256,512)$ for the cases with $N_{\rho}=3.0$, and $(N_{r},N_{\theta},N_{\phi})=(241,320,640)$ for the cases with $N_{\rho}=4.6$. In order to explore how sensitive the magnitudes in Fig. 9 are, we vary the Prandtl, magnetic Prandtl, and Rayleigh numbers, the value of $N_{\rho}$, and the boundary conditions. Generally speaking, the models shown above reach a steady state dynamo with equatorial jets. More details and the main results of these simulations will be presented in an upcoming paper dedicated to the dynamo in HJs, closely following the methodology employed in Elias-López et al. (2025a, b). For the specific purpose of this work, we are only interested in the profiles shown in Fig. 9.
@@ Appendix C Results from standard cooling models
Figure 13: Evolution of irradiated HJs with no additional heating, $\epsilon_{j}=0$, for models with $M=1\leavevmode\nobreak\ M_{j}$, $\rho_{c}=10$  g cm-3 (top row) or $M=4\leavevmode\nobreak\ M_{j}$, $\rho_{c}=45$  g cm-3 (bottom row). Left panels: $p(T)$ profiles at ages tsimilar-$t\sim$ 0.1, 1 and 5 Gyr (dots, dashes, solid lines, respectively), comparing the cases without (black) and with irradiation, $T_{\rm eq}=1500$ K (blue) and 2000 K (orange), all with solar composition. The thicker lines correspond to the convective zone, which extends from the RCB ($p\sim 1-10^{3}$ bar) to the outer core ($p>10^{7}$ bar). Central panels: Radius evolution of planets with $T_{\rm eq}=1500$ K and different core masses ($M_{c}=0,10\leavevmode\nobreak\ (M/M_{j}),50\leavevmode\nobreak\ (M/M_{j})\leavevmode\nobreak\ M_{\oplus}$, increasing for darker shades, all with solar composition) and composition of the envelope: pure hydrogen (purple), solar (black) and higher helium and heavier elements content, $(Y,Z)=(0.35,0.035)$ (purple, black and red, respectively, all with $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$). In the top central panel we also show the effects of variation of $\rho_{c}=5$ g cm-3 (dashes) and $\rho_{c}=20$ g cm-3 (dot-dashed lines) for the representative $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$, solar composition case: since the effect is rather small compared to varying other parameters, we show it only in that particular panel. Right panels: Same as the central panels, but for $T_{\rm eq}=2000$ K.
@@
As a reference and comparison with well-known results of classical cooling, we briefly show illustrative examples for the planetary structure with no extra heating, $\epsilon_{j}=0$ in eq. (5). In Fig. 13 we summarize the main results for a few representative cases with $M=1\leavevmode\nobreak\ M_{j}$ (top panels) and $M=4\leavevmode\nobreak\ M_{j}$ (bottom panels). In the left panels we compare the $p(T)$ profiles of the non-irradiated case (black), and HJs with $T_{\rm eq}=1500$ K (blue) and 2000 K (orange), at three different ages: 0.1, 1 and 5 Gyr (different line styles). In the same plot, we indicate the position of the convective (thick lines) and radiative layers (thin lines). In all cases, as the internal heat decreases over time, the $p(T)$ lines naturally move downwards. For a cold Jupiter, the radiative-convective boundary (RCB) always remains very shallow ($\sim 1$ bar), and additional radiative regions at intermediate pressure can appear as the internal entropy drops (Ginzburg & Sari 2015; Komacek & Youdin 2017). For HJs, instead, irradiation causes much thicker outermost layers to become radiative. The specific profiles depend on the atmospheric modelling: in the simple MESA recipes, the profiles are isothermal, with a drop in temperature in the region $m<\Sigma_{\star}$; non-grey atmospheric models provide non-isothermal, longitude-dependent profiles, including thermal inversion, e.g., Gandhi & Madhusudhan (2019). Such details are of primary importance for atmospheric modelling and data interpretation, but are not relevant for this work, since HJ radius inflation has to be driven from much deeper layers.
@@
Regardless of these details, the irradiation pushes the RCB to much deeper levels, typically $p\sim 10^{3}$ bar at Gyr ages. Moreover, when comparing the $p(T)$ curves of different $T_{\rm eq}$, the blanketing effect of the highly irradiated layers is evident: the interior is kept warmer for HJs with higher $T_{\rm eq}$. These preliminary considerations on irradiation and RCB are in agreement with what has been widely discussed in previous works, e.g. Arras & Bildsten (2006); Fortney et al. (2007); Ginzburg & Sari (2015); Komacek & Youdin (2017); Thorngren et al. (2019).
@@
In the central panels of Fig. 13, we show how the irradiation (the same cases as the left panel), core mass and density impact the radius evolution. We vary the core mass: $M_{c}=(0,10,50)\leavevmode\nobreak\ M_{\oplus}$. For the composition, we vary the baseline values (solar), considering two additional cases: pure hydrogen, $Y=Z=0$, and heavy composition, with $(Y,Z)=(0.35,0.035)$, covering the upper range of the dispersion observed in stellar helium abundances and metallicities (Verma et al. 2019). Higher values of metallicities are not discarded (Thorngren et al. 2016), and have the effect of further contracting the planet. Since we are focusing on explaining the large inflated radii, we don’t explore them further. We also show the negligible effect of varying the core density alone by a factor of 2 (dashed and dot-dashed lines in the top central panel): in the rest of the paper, we will use fixed, standard values of $\rho_{c}(M)$, as presented above.
@@
On one side, as expected, we find positive trends between radius inflation and irradiation. The radius can increase up to $\sim 20\%$ at Gyr ages for $T_{\rm eq}=2000$ K, for $M=1\leavevmode\nobreak\ M_{j}$, $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$, $\rho_{c}=10$  g cm-3, and solar composition. On the other hand, higher values of $M_{c}$ or $\rho_{c}$ make the planet more compact, while higher metallicity makes the planets cool faster. When we look at heavier planets, of which we show the $M=4\leavevmode\nobreak\ M_{j}$ cases in the bottom rows of Fig. 13, the dispersion in the cooling curve using the same ranges of parameters is similar to the $M=1\leavevmode\nobreak\ M_{j}$ HJs. This is due to the total higher inertia (and the lower relative weight of the core for the range considered here).
@@
In summary, the total spread of $R(t)$ for HJs at Gyr ages is in the range of $\sim 1-1.4\leavevmode\nobreak\ R_{j}$, with typical values around $1.2\leavevmode\nobreak\ R_{j}$. This is due to the combined effects of, in order of importance, $T_{\rm eq}$, $M$, $M_{c}$, and composition, and is compatible with earlier results by e.g. Guillot et al. (1996); Fortney et al. (2007); Thorngren & Fortney (2018). These results represent the baseline cooling model, on top of which we add the Ohmic heating model, as described next.
@@
In Fig. 14 we show the sensitivity of the models on a range of values $\Sigma_{\star}\in[100,500]$  g cm-2, in case of $T_{\rm eq}=1500$ and 2000 K. This value approximately corresponding to homogeneous (per unit mass) absorption of the irradiation down to pressures from $\sim$0.5 bar to $\sim$1.2 bar for Gyr ages. Remember that $\Sigma_{\star}$ effectively models in a simple way how deeply the radiation is absorbed, which in reality depends on the specific stellar spectrum and the wavelength-dependent opacities (i.e., composition and atmospheric structure) of each planet. We see an important effect for early ages, but at Gyr ages, the impact is few percent at most. Note that Owen & Wu (2016), in the context of atmospheric mass loss by highly irradiated super Earths, proved a similar limited sensitivity of results on a broader range $\Sigma_{\star}\in[50,1250]$  g cm-2.
@@
Figure 14: Evolution of the radius for models with $M=1\leavevmode\nobreak\ M_{j}$, $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$, $\rho_{c}=10$ g cm-3, solar composition, no Joule heating, comparing cases with $T_{\rm eq}=1500$ K (blue) and 2000 K (orange), and $\Sigma_{\star}=100,250,500$  g cm-2 (dashes, solid lines, dots, respectively).

### v2

@@ B.1 Jovian magnetic field curvature
where we have separated the internal (dynamo-generated) and external (magnetosphere/magneto-disk) components. Public data is given in terms of the set of constants $g_{n}^{m}$, $h_{n}^{m}$, $G_{n}^{m}$ and $H_{n}^{m}$ (the Schmidt coefficients) which define the spherical harmonic expansions up to some specific multipole degree using the Schmidt quasi-normalised associated Legendre polynomials $P_{n}^{m}(\cos\theta)$,
@@
Figure 12: Inverse of the mean curvature modulus for different Jupiter magnetic field models as a function of radius, obtained using a discretised version of equation (26). The JRM33 models do not significantly change if the external multipoles are included.
Fig. 12 shows the inverse of the mean curvature at different radii for both models, including two different cut-offs for JRM33. The gray zone approximately marks the possible start of the conductive zone where the gradual appearance of metallic hydrogen would invalidate the potential solution. The inverse curvature has a mean value of the order of $1/\kappa\sim 0.1\leavevmode\nobreak\ R_{\rm j}$ at $r\sim(0.8-0.9)\leavevmode\nobreak\ R_{\rm j}$ (where the dynamo surface should be), which can be taken as a lower limit of the magnetic field curvature inside the Jovian dynamo.
@@ B.2 Details of MHD dynamo simulations
For example, for the models shown in Fig. 9, the background MESA profiles correspond to an irradiated $1M_{\rm j}$ planet with internal Ohmic heating at 10 Gyr. At that time, the simulated planetary radius is $1.218\leavevmode\nobreak\ R_{\rm j}$, corresponding to an external pressure of and 64 and 5.4 kbar, respectively. We used resolutions of $(N_{r},N_{\theta},N_{\phi})=(193,256,512)$ for the cases with $N_{\rho}=3.0$, and $(N_{r},N_{\theta},N_{\phi})=(241,320,640)$ for the cases with $N_{\rho}=4.6$. In order to explore how sensitive the magnitudes in Fig. 9 are, we vary the Prandtl, magnetic Prandtl, and Rayleigh numbers, the value of $N_{\rho}$, and the boundary conditions. Generally speaking, the models shown above reach a steady state dynamo with equatorial jets. More details and the main results of these simulations will be presented in an upcoming paper dedicated to the dynamo in HJs, closely following the methodology employed in Elias-López et al. (2025a, b). For the specific purpose of this work, we are only interested in the profiles shown in Fig. 9.
@@ Appendix C Results from standard cooling models
Figure 13: Evolution of irradiated HJs with no additional heating, $\epsilon_{\rm j}=0$, for models with $M=1\leavevmode\nobreak\ M_{\rm j}$, $\rho_{c}=10$  g cm-3 (top row) or $M=4\leavevmode\nobreak\ M_{\rm j}$, $\rho_{c}=45$  g cm-3 (bottom row). Left panels: $p(T)$ profiles at ages tsimilar-$t\sim$ 0.1, 1 and 5 Gyr (dots, dashes, solid lines, respectively), comparing the cases without (black) and with irradiation, $T_{\rm eq}=1500$ K (blue) and 2000 K (orange), all with solar composition. The thicker lines correspond to the convective zone, which extends from the RCB ($p\sim 1-10^{3}$ bar) to the outer core ($p>10^{7}$ bar). Central panels: Radius evolution of planets with $T_{\rm eq}=1500$ K and different core masses ($M_{c}=0,10\leavevmode\nobreak\ (M/M_{\rm j}),50\leavevmode\nobreak\ (M/M_{\rm j})\leavevmode\nobreak\ M_{\oplus}$, increasing for darker shades, all with solar composition) and composition of the envelope: pure hydrogen (purple), solar (black) and higher helium and heavier elements content, $(Y,Z)=(0.35,0.035)$ (purple, black and red, respectively, all with $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$). In the top central panel we also show the effects of variation of $\rho_{c}=5$ g cm-3 (dashes) and $\rho_{c}=20$ g cm-3 (dot-dashed lines) for the representative $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$, solar composition case: since the effect is rather small compared to varying other parameters, we show it only in that particular panel. Right panels: Same as the central panels, but for $T_{\rm eq}=2000$ K.
@@
As a reference and comparison with well-known results of classical cooling, we briefly show illustrative examples for the planetary structure with no extra heating, $\epsilon_{\rm j}=0$ in eq. (5). In Fig. 13 we summarise the main results for a few representative cases with $M=1\leavevmode\nobreak\ M_{\rm j}$ (top panels) and $M=4\leavevmode\nobreak\ M_{\rm j}$ (bottom panels). In the left panels we compare the $p(T)$ profiles of the non-irradiated case (black), and HJs with $T_{\rm eq}=1500$ K (blue) and 2000 K (orange), at three different ages: 0.1, 1 and 5 Gyr (different line styles). In the same plot, we indicate the position of the convective (thick lines) and radiative layers (thin lines). In all cases, as the internal heat decreases over time, the $p(T)$ lines naturally move downwards. For a cold Jupiter, the radiative-convective boundary (RCB) always remains very shallow ($\sim 1$ bar), and additional radiative regions at intermediate pressure can appear as the internal entropy drops (Ginzburg & Sari 2015; Komacek & Youdin 2017). For HJs, instead, irradiation causes much thicker outermost layers to become radiative. The specific profiles depend on the atmospheric modelling: in the simple MESA recipes, the profiles are isothermal, with a drop in temperature in the region $m<\Sigma_{\star}$; non-grey atmospheric models provide non-isothermal, longitude-dependent profiles, including thermal inversion, for example, Gandhi & Madhusudhan (2019). Such details are of primary importance for atmospheric modelling and data interpretation, but are not relevant for this work, since HJ radius inflation has to be driven from much deeper layers.
@@
Regardless of these details, the irradiation pushes the RCB to much deeper levels, typically $p\sim 10^{3}$ bar at gigayear ages. Moreover, when comparing the $p(T)$ curves of different $T_{\rm eq}$, the blanketing effect of the highly irradiated layers is evident: the interior is kept warmer for HJs with higher $T_{\rm eq}$. These preliminary considerations on irradiation and RCB are in agreement with what has been widely discussed in previous works, e.g. Arras & Bildsten (2006); Fortney et al. (2007); Ginzburg & Sari (2015); Komacek & Youdin (2017); Thorngren et al. (2019).
@@
In the central panels of Fig. 13, we show how the irradiation (the same cases as the left panel), core mass and density impact the radius evolution. We vary the core mass: $M_{c}=(0,10,50)\leavevmode\nobreak\ M_{\oplus}$. For the composition, we vary the baseline values (solar), considering two additional cases: pure hydrogen, $Y=Z=0$, and heavy composition, with $(Y,Z)=(0.35,0.035)$, covering the upper range of the dispersion observed in stellar helium abundances and metallicities (Verma et al. 2019). Higher values of metallicities are not discarded (Thorngren et al. 2016), and have the effect of further contracting the planet. Since we are focusing on explaining the large inflated radii, we do not explore them further. We also show the negligible effect of varying the core density alone by a factor of 2 (dashed and dot-dashed lines in the top central panel): In the rest of the paper, we use fixed standard values of $\rho_{c}(M)$, as presented above.
@@
On one side, as expected, we find positive trends between radius inflation and irradiation. The radius can increase up to $\sim 20\%$ at gigayear ages for $T_{\rm eq}=2000$ K, for $M=1\leavevmode\nobreak\ M_{\rm j}$, $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$, $\rho_{c}=10$  g cm-3, and solar composition. On the other hand, higher values of $M_{c}$ or $\rho_{c}$ make the planet more compact, while higher metallicity makes the planets cool faster. When we look at heavier planets, of which we show the $M=4\leavevmode\nobreak\ M_{\rm j}$ cases in the bottom rows of Fig. 13, the dispersion in the cooling curve using the same ranges of parameters is similar to the $M=1\leavevmode\nobreak\ M_{\rm j}$ HJs. This is due to the total higher inertia (and the lower relative weight of the core for the range considered here).
@@
In summary, the total spread of $R(t)$ for HJs at gigayear ages is in the range of $\sim 1-1.4\leavevmode\nobreak\ R_{\rm j}$, with typical values around $1.2\leavevmode\nobreak\ R_{\rm j}$. This is due to the combined effects of, in order of importance, $T_{\rm eq}$, $M$, $M_{c}$, and composition, and is compatible with earlier results by e.g. Guillot et al. (1996); Fortney et al. (2007); Thorngren & Fortney (2018). These results represent the baseline cooling model, on top of which we add the Ohmic heating model, as described next.
@@
In Fig. 14 we show the sensitivity of the models on a range of values $\Sigma_{\star}\in[100,500]$  g cm-2, in case of $T_{\rm eq}=1500$ and 2000 K. This value approximately corresponding to homogeneous (per unit mass) absorption of the irradiation down to pressures from $\sim$0.5 bar to $\sim$1.2 bar for gigayear ages. Remember that $\Sigma_{\star}$ effectively models in a simple way how deeply the radiation is absorbed, which in reality depends on the specific stellar spectrum and the wavelength-dependent opacities (i.e. composition and atmospheric structure) of each planet. We see an important effect for early ages, but at gigayear ages, the impact is of a few percent at most. We note that Owen & Wu (2016), in the context of atmospheric mass loss by highly irradiated super Earths, proved a similar limited sensitivity of results on a broader range $\Sigma_{\star}\in[50,1250]$  g cm-2.
@@
Figure 14: Evolution of the radius for models with $M=1\leavevmode\nobreak\ M_{\rm j}$, $M_{c}=10\leavevmode\nobreak\ M_{\oplus}$, $\rho_{c}=10$ g cm-3, solar composition, no Joule heating, comparing cases with $T_{\rm eq}=1500$ K (blue) and 2000 K (orange), and $\Sigma_{\star}=100,250,500$  g cm-2 (dashes, solid lines, dots, respectively).